OpenStudy (anonymous):

The price (in dollars) p and the quantity demanded q are related by the equation: p^2+2q^2=1100.If R is revenue, dR/dt can be expressed by the following equation: dR/dt=A*dp/dt, where A is a function of just q. A = ? Find dR/dt when q=15 and dp/dt=5. dR/dt = The price (in dollars) p and the quantity demanded q are related by the equation: p^2+2q^2=1100.If R is revenue, dR/dt can be expressed by the following equation: dR/dt=A*dp/dt, where A is a function of just q. A = ? Find dR/dt when q=15 and dp/dt=5. dR/dt = @Mathematics

myininaya (myininaya):

\[2pp'+4qq'=0 =>p'=\frac{-2 q q'}{p} => R'=Ap'=A[\frac{-2 q q'}{p}]\] So I guess you can solve this for A \[R'=A [\frac{-2 q q'}{p} ]\] I'm not sure if that is what you want since you said A=f(q)